OpenStudy (anonymous):

PLEASE HELP ME!! 4x-12 over 3x-9

OpenStudy (shamil98):

4x-12/3x-9 = 4(x-3)/3(x-3) = 4/3

OpenStudy (anonymous):

THANK YOU! you are soo helpful :) thanks for showing ur work it was a referesher!

OpenStudy (shamil98):

np :)